check continuity and differentiability


f(x)= {1/1+ e^1/x} , x is not equal to 0


         0   , x=0

Grade:12

1 Answers

Jitender Singh IIT Delhi
9 years ago
Ans:
Checking the continuity:
Left hand limit(LHL) = Right hand limit(RHL) = f(0)
LHL = \lim_{h\rightarrow 0}\frac{1}{1+e^{\frac{-1}{h}}}
LHL = \lim_{h\rightarrow 0}\frac{e^{\frac{1}{h}}}{1+e^{\frac{1}{h}}}
LHL = \lim_{h\rightarrow 0}\frac{e^{\frac{1}{h}}+1-1}{1+e^{\frac{1}{h}}}
LHL = \lim_{h\rightarrow 0}1-\frac{1}{1+e^{\frac{1}{h}}} = 1
RHL = \lim_{h\rightarrow 0}\frac{1}{1+e^{\frac{1}{h}}} = 0
f(0) = 0
So, f(x) is not continuous at x = 0.
Similarly, you can check the left hand & right hand derivative.
\lim_{h\rightarrow 0}\frac{f(-h)}{-h} = \lim_{h\rightarrow 0}\frac{f(h)}{h}
Thanks & Regards
Jitender Singh
IIT Delhi
askIITians Faculty